OpenStudy (anonymous):

suppose that the population of deer in state is 1,500 and is growing 2% each year.predicit the population after 4 years. a. about 12,000 deer. b.about 1,624 deer. c. about 3,110 deer. d. about 24,000 deer. need help.and explain plz.

OpenStudy (anonymous):

\[P_n=1500·(1+0.02)^n=1500·1.02^n\] Solve for n=4

OpenStudy (anonymous):

The formula is obtained like this: Current Population=P0=1500 After 1 year: P1=P0+0.02·P0=P0·(1+0.02) After 2 years:P2=P1+0.02·P1=P1·(1+0.02)=P0·(1+0.02)·(1+0.02)=P0·(1+0.02)^2 After 3 Years:P3=P2·(1+0.02)=P0·(1+0.02)^2(1+0.02)=P0·(1+0.02)^3 Same rationale to get P4=P0·(1+0.02)^4=1500·1.02^4

OpenStudy (anonymous):

so the answer will be a.12,000 deer

OpenStudy (anonymous):

wrong calculation, try again

OpenStudy (anonymous):

Give me 1.02^4

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

give me 1.02^4

OpenStudy (anonymous):

1.08

OpenStudy (anonymous):

Now 1.08 times 1500

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

1.02^4 is not 1.08

OpenStudy (anonymous):

got it the answer is B. 1,624 deer

OpenStudy (anonymous):

It is a bit more than 1.08

OpenStudy (anonymous):

@redapple then the answer is:????

OpenStudy (anonymous):

do you think is 3,110 deer

OpenStudy (anonymous):

It is B

OpenStudy (anonymous):

thanx 4 the help.

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

(9*10^5)(6*10^-7) a. 5.4*10^-34 b.1.5*10^-1 c.5.4*10^-1 d.1.5*10^34 need help plz.

OpenStudy (anonymous):

Some help for you:\[(x·10^n)·(y·10^m)=(x·y)·10^{(n+m)}\]